The Golden Eagle Corporation has the following items on their income and balance sheets (values in tables are in thousands: Balance Sheet Items Assets) Last Year Two Years Ago 520 500 Cash Accounts Receivable 530 540 Inventory 286 286 Total Current Assets Fixed Assets 4960 45 2096 Depreciation Net Fixed Assets Balance Sheet Items Liabilities)
media%2F552%2F552a5c0a-92b9-4fa4-b23a-25
media%2Fa6a%2Fa6afba44-7980-4e62-819f-44
Answer #1
ROE = Profit Margin x Asset Turnover x Equity Multiplier
Using Average Assets to calculate Assets Turnover
Profit Margin = Net income - preferred / Sales
Net Income = $3080 - 1870 - 560 - 94 -85 -96 $375.00
Profit Margin = $375/$3080 12.18%
Asset Turnover = Sales/ Average Total Assets
Asset Turnover = $3080 /((3991+4190)/2) 0.75 times
Total Assets past year = 520+530+286+5000-2345 3991
Total Assets 2 years ago = 500+540+286+4960-2096 4190
Equity Multiplier = Total assets/ Total Equity
Equity Multiplier = 3991 /(340+1390) 2.3
ROE = 12.18% x .75 x 2.3 21.15%
ROE = Net income available to common stockholders/ Shareholder Equity
21.15% = Net income available to common stockholders/(340+1390) $365.88
Using Total Assets to calculate Asset Turnover
Profit Margin = Net income - preferred / Sales
Net Income = $3080 - 1870 - 560 - 94 -85 -96 $375.00
Profit Margin = $375/$3080 12.18%
Asset Turnover = Sales/ Average Total Assets
Asset Turnover = $3080 /3991 0.77 times
Total Assets past year = 520+530+286+5000-2345 3991
Equity Multiplier = Total assets/ Total Equity
Equity Multiplier = 3991 /(340+1390) 2.3
ROE = 12.18% x .75 x 2.3 21.68%
ROE = Net income available to common stockholders/ Shareholder Equity
21.15% = Net income available to common stockholders/(340+1390) $375.00